OpenStudy (anonymous):

Simplify the following expression x - 4y + 5z - 3x - y + 2z = ?

pooja195 (pooja195):

ok lets combine the like terms whats -3x+x ?

OpenStudy (mathmath333):

this is same as \(\large \color{black}{\begin{align} x - 4y + 5z - 3x - y + 2z &=x- 3x - y- 4y + 2z+ 5z \hspace{.33em}\\~\\ &=x(1- 3) - y(1+4) + z(2+ 5) \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

-2x @pooja195

OpenStudy (mathmath333):

yep do it for \(-4y-y \) and \(5z+2z\) also

OpenStudy (anonymous):

-2x -3y +7x

OpenStudy (mathmath333):

yes \(\Huge \color{black}{\begin{align} \checkmark \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

your answer is correct